OpenStudy (anonymous):

EASY! MEDAL AND FAN! What is the kinetic energy of a 45 kg object moving at 13m/s

OpenStudy (isaiah.feynman):

\[K = \frac{ 1 }{ 2 }m v^{2}\]

OpenStudy (anonymous):

What represents m?

OpenStudy (isaiah.feynman):

You have m and v given, substitute.

OpenStudy (isaiah.feynman):

M is mass. Mass is measured in kilograms.

OpenStudy (anonymous):

So it would be K=1/245x13^2?

OpenStudy (anonymous):

Im very bad at this, Appreciate the help. :)

OpenStudy (isaiah.feynman):

Yep.

OpenStudy (anonymous):

Can you break it down real quick on how to do this? Im sorry, Trying to help my son...

OpenStudy (isaiah.feynman):

You're on the right track. Put 13 for for v and 45 for m.

OpenStudy (anonymous):

Yes, Then what? Do I square 13?

OpenStudy (isaiah.feynman):

Yes. Like this \[K= \frac{ 1 }{ 2 }(45)(13)^{2}\]

OpenStudy (isaiah.feynman):

\[13^{2} = 169\]So we can just do \[K = \frac{ 1 }{ 2 }(45)(169)\]

OpenStudy (anonymous):

169

OpenStudy (isaiah.feynman):

Just multiply through.

OpenStudy (anonymous):

45x169?

OpenStudy (isaiah.feynman):

Yes. times 1/2

OpenStudy (anonymous):

7605 / 2?

OpenStudy (isaiah.feynman):

Yes.

OpenStudy (isaiah.feynman):

Then you get your answer. I have to go to work. :)

OpenStudy (anonymous):

3802.5